OpenStudy (mysticfury):

The measure of an angle is 63°. What is the measure of its complement? A. 27° B. 37° C. 63° D. 117° C or D?

OpenStudy (mysticfury):

@LightLegendX @annah.bee

OpenStudy (mysticfury):

@ashutoshrcf

OpenStudy (babeberry):

90-63=?

TheSmartOne (thesmartone):

complementary angles add up to 90 degrees 63 + x = 90 x = 90 - 63 x = ??

OpenStudy (mysticfury):

oh okay thanks guys!

TheSmartOne (thesmartone):

Anytime! :)